# Aleph-Alpha-GermanWeb: Improving German-language LLM pre-training with model-based data curation and synthetic data generation

### v1

@@ 1 Introduction
    removing semantic duplication, e.g., otherwise identical documents with slightly different wording, which are typically straightforward to identify and remove [24];
@@
    removing semantic redundancy, e.g., different summaries of the same book, which we expect to maintain some of for the purposes of syntactic variance, but should probably only keep in proportion to the level of importance or complexity of the underlying concept [1];
@@
    augmenting or upsampling good data, e.g., generating synthetic data using LLMs [27] or intentionally duplicating high-quality subsets of the data [21].
@@
In this paper we present a data curation pipeline specifically for German-language LLM pre-training. We use it to derive our own dataset, Aleph-Alpha-GermanWeb, from three sources: (1) Common Crawl [13] web data not included in FineWeb2; (2) FineWeb2; and (3) synthetic data conditioned on actual, organic web data. Our contributions are:
@@ 2 Methods
Our data curation methodology consists of three approaches: a data filtering pipeline for Common Crawl documents, synthetic document generation, and quality bucketing of all documents.
@@ 2.1 Curating Common Crawl
We followed the RefinedWeb pipeline [39], as adapted for German in FineWeb2 [21], with some adjustments. We implemented all steps of the pipeline using NeMo Curator [23]. We summarise the number of web documents our pipeline yielded in Table 1.
@@
We downloaded six Common Crawl [13] web data dumps, spanning from September 2024 to February 2025. Each dump captured raw web page data, e.g., HTML, metadata such as URLs, and text extracts.
@@ 2.2 Synthetic data generation
Inspired by similar work for English-language web data [28, 46], we performed synthetic data generation conditioned on actual – what we also refer to as ‘organic’ – data from FineWeb2. We showcase the five prompt templates we used for synthetic data generation in Appendix B, which are designed to generate educational or basic rephrasings, summarizations, question-answer pairs, and lists of factual information. For all synthetic generation, we used the 12 billion parameter Mistral-Nemo-Instruct-2407111https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 LLM.
@@ 2.3 Quality classification
Following prior work on quality classification for English-language web data [46], we categorised documents into five quality buckets – high, medium-high, medium, medium-low, and low – by leveraging the strengths of several imperfect quality classifiers to produce a more accurate outcome. Specifically, we created several individual quality classifiers based on fastText [7] and Bidirectional Encoder Representations from Transformers (BERT) [11] models. In the following, we describe (1) our two grammar classifiers, (2) our two educational quality classifiers, and (3) the way in which we ensembled all classifiers and grouped the data into the five quality buckets according to the classifiers’ scores.
@@ 2.3.2 Educational quality
Inspired by successful demonstrations for English with FineWeb [38] and Nemotron-CC [46], we created educational quality classifiers to curate high-quality German web data by training a classifier on scores given by an LLM-as-a-judge. The intuition behind this is to identify web documents that are more likely to be educational, informative, and useful for model training, as opposed to spam, low-quality, or otherwise unhelpful content.
@@
For the binary fastText classifier, the low- and high-quality training data subsets consisted of 185,403 documents each. The low quality subset consisted of documents with educational quality scores of one or two, whereas the high quality subset consisted of documents scoring four or five. We used 95% of the data (and the remaining 5% for validation) to train a fastText model to classify between high- and low-quality data. It reached 92% precision and 91.5% recall on the validation set.
@@ 2.3.3 Ensembling
Model Condition Points Initialise all documents 0 BERT (edu. quality) Predicted score = 5 +3 BERT (edu. quality) Predicted score = 4 +2 BERT (edu. quality) Predicted score = 3 +1 fastText (edu. quality) Predicted “high quality” with >99% confidence +3 fastText (edu. quality) Predicted “high quality” with >95% and $\leq$99confidence +2 BERT (grammar) Predicted “high quality” +2 fastText (grammar) Predicted “high quality” with >99% confidence +3 fastText (grammar) Predicted “high quality” with >95% and $\leq$99confidence +2 Document < 100 characters and score > 6 –3
@@
Quality Bucket Score Range Quality Distribution
Filtered CC Synthetic FineWeb2
Docs Bytes Docs Bytes Docs Bytes
High $\geq 10$ 20.7% 10.5% 22.1% 17.9% 16.5% 6.9%
Medium-high $8-9$ 23.7% 18.4% 28.0% 26.4% 19.5% 12.1%
Medium $6-7$ 17.3% 16.7% 21.9% 23.6% 15.4% 11.9%
Medium-low $4-5$ 14.8% 17.2% 14.6% 15.9% 14.8% 14.3%
Low $<4$ 23.4% 37.2% 13.4% 16.3% 33.8% 54.8%
@@ 3.1 Evaluations
Figure 1: Average accuracies (MMMLU, ARC-Easy, HellaSwag; single- and five-shot) of 1B Llama-style models trained on $\sim$ 84 billion tokens from different datasets. Our combined dataset, Aleph-Alpha-GermanWeb, and synthetic subset alone, outperforms FineWeb2. Here, 1,000 training steps equates to approximately 2.1 billion tokens. Individual benchmark results are provided in Appendix D.1.
@@
We trained and evaluated our filtered Common Crawl (Filtered CC), synthetic, and an equal mixture of Filtered CC, synthetic, and a subset of FineWeb2 (1-1-1 Mix), and compared these to FineWeb2. The results are shown in Figure 1. We chose to exclude TruthfulQA from the 1B parameter model evaluations, as we found that it did not give a clear signal and suspect that this size of the model may be incapable of the task (cf. §3.3 for results of an 8B model).
@@
Figure 2 compares the performance of these models pre-trained for different datasets and training scenarios. Differently to the 1B Llama-style models, these 8B HAT models operate on words, not tokens, resulting in 1,000 training steps equating to approximately three billion words in this set-up. We confirmed our previous finding of our filtered CC data outperforming FineWeb2 at the 1B scale in Figure 2 (left). We also confirmed that our synthetic data on average outperforms randomly-sampled FineWeb2 data in Figure 2 (centre).
When pre-training LLMs, web data is often mixed from different languages and with data from alternative, curated sources [16], e.g., English and German data might be mixed with additional non-web data from books and encyclopedias. To test the performance of our single highest-performing subset, we compared our synthetic data to mix of 50% FineWeb2 and 50% high-quality curated datasets [^4] In particular: the German National Library, European Union Parliament transcripts, European Patent Office applications, German Text Archive, Project Gutenberg books, JRC-Acquis, NewsAPI, Open Legal Data, Open Web Math, peS2o, PhilPapers, Wikibooks, Wikinews, Wikpedia, Wikisource, and Wikivoyage.. Figure 2 (right) shows our synthetic data outperformed this mixture on our tested benchmarks.
Figure 2: Accuracies of 8B HAT models trained on different datasets. On average, and for all but one individual benchmark comparison, Aleph-Alpha-GermanWeb outperforms FineWeb2. (Left) After 50,000 training steps, equating to $\sim$ 150 billion words. (Centre) After 21,000 training steps, equating to $\sim$ 63 billion words. (Right) After 25,000 training steps on an English-language web-derived dataset, followed by 20,000 training steps on the respective German-language datasets, equating to $\sim$ 75 billion English words followed by $\sim$ 60 billion German words.
@@ 4 Discussion and Conclusions
Data preparation: Michael Barlow, Bastian Harren, Volker Stampa, Gregor Ziegltrum
@@ Appendix D Extended results
Overall, subsets of Aleph-Alpha-GermanWeb and their combination outperform FineWeb2 across all benchmarks. Interestingly, we find that different subsets have different strengths: our synthetic and high-quality classified data performed particularly well on MMMLU, whereas our filtered CC data performed best for HellaSwag.

### v2

@@ 1 Introduction
    removing syntactic (near) duplication, e.g., identical documents (or with slightly different wording), which are typically straightforward to identify and remove [24];
@@
    removing semantic duplication, e.g., different summaries of the same book, which we expect to maintain some of for the purposes of syntactic variance, but should probably only keep in proportion to the level of importance or complexity of the underlying concept [1];
@@
    augmenting or upsampling good data, e.g., adding high-quality synthetic data generated by LLMs [27] or intentionally duplicating high-quality subsets of the data [21].
@@
In this paper we present a data curation pipeline specifically for German-language LLM pre-training. We use it to derive our own dataset, Aleph-Alpha-GermanWeb, from three sources: (1) Common Crawl [13] web data not included in FineWeb2; (2) FineWeb2 (ODC-By v1.0); and (3) synthetic data conditioned on actual, organic web data. Our contributions are:
@@ 2 Methods
Our data curation methodology consists of three approaches: a data filtering pipeline for Common Crawl documents, synthetic document generation, and quality bucketing of all FineWeb2 documents.
@@ 2.1 Curating Common Crawl
We followed the RefinedWeb pipeline [39], as adapted for German in FineWeb2 [21], with some adjustments. We implemented all steps of the pipeline using NeMo Curator [23] (Apache-2.0). We summarise the number of web documents our pipeline yielded in Table 1.
@@
We downloaded six Common Crawl [13] (CC BY-SA 4.0) web data dumps, spanning from September 2024 to February 2025. Each dump captured raw web page data, e.g., HTML, metadata such as URLs, and text extracts.
@@ 2.2 Synthetic data generation
Inspired by similar work for English-language web data [28, 46], we performed synthetic data generation conditioned on actual – what we also refer to as ‘organic’ – data from FineWeb2. We showcase the five prompt templates we used for synthetic data generation in Appendix B, which are designed to generate educational or basic rephrasings, summarizations, question-answer pairs, and lists of factual information. For all synthetic generation, we used the 12 billion parameter Mistral-Nemo-Instruct-2407111https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 (apache 2.0) LLM.
@@ 2.3 Quality classification
Following prior work on quality classification for English-language web data [46], we categorised documents into five quality buckets – high, medium-high, medium, medium-low, and low – by leveraging the strengths of several imperfect quality classifiers to produce a more accurate outcome. Specifically, we created several individual quality classifiers based on fastText [7] and Bidirectional Encoder Representations from Transformers (BERT) [11] models. In the following, we describe (1) our two grammar classifiers, (2) our two educational quality classifiers, (3) our two instruction classifiers and (4) the way in which we ensembled all classifiers and grouped the data into the five quality buckets according to the classifiers’ scores.
@@ 2.3.2 Educational quality
Inspired by successful demonstrations for English with FineWeb [38] and Nemotron-CC [46] (Apachee 2.0), we created educational quality classifiers to curate high-quality German web data by training a classifier on scores given by an LLM-as-a-judge. The intuition behind this is to identify web documents that are more likely to be educational, informative, and useful for model training, as opposed to spam, low-quality, or otherwise unhelpful content.
@@
For the binary fastText classifier, the low- and high-quality training data subsets consisted of 185,403 documents each. The low quality subset consisted of documents with educational quality scores of one or two, whereas the high quality subset consisted of documents scoring four or five. We used 95% of the data (and the remaining 5% for validation) to train a fastText model to classify between high- and low-quality data. It reached 77% precision and 77% recall on the validation set.
@@
2.3.3 Instruction
Building on Messmer et al. [31], we trained binary classifiers using both fastText and BERT to distinguish between chat/instruction-style data (high quality) and other types of documents (low quality). To create a balanced training dataset, we created two subsets of 80,000 documents each. The low-quality subset consisted of randomly selected documents from the German FineWeb2 corpus, while the high-quality subset comprised randomly sampled instances from the Aya collection, Include Base-44, OpenAssistant2, and MMLU. To create samples from MMLU, we first concatenated a question and the respective answers. For OpenAssistant-2, we concatenated all samples that belong to the same conversation. We used 95% of the data for training and the remaining 5% for validation.
The fastText classifier indicated strong discriminative performance, achieving 99% precision and 99% recall on the validation set. Similarly, the BERT-based classifier reached a validation accuracy of 100%.
@@ 2.3.4 Ensembling
Model Condition Points Initialise all documents 0 BERT (edu. quality) Predicted score = 5 +3 fastText (edu. quality) Predicted “high quality” with >99% confidence +2 BERT (grammar) Predicted “high quality” +3 fastText (grammar) Predicted “high quality” with >99% confidence +2 BERT (instruction) High quality probability is in top 15% of the distribution +6 fastText (instruction) High quality probability is in top 15% of the distribution +4
@@
Quality Bucket Score Range Quality Distribution of FineWeb2
Docs Bytes
High $\geq 12$ 12.1% 8.1%
Medium-high $9-11$ 15.0% 9.9%
Medium $5-8$ 36.3% 28.7%
Medium-low $3-4$ 15.5% 21.7%
Low $<3$ 21.0% 31.6%
@@ 3.1 Evaluations
Figure 1: Average accuracies (MMMLU, ARC-Easy, HellaSwag; single- & five-shot) of 1B Llama-style models trained on $\sim$ 84 billion tokens from different datasets. All subsets of Aleph-Alpha-GermanWeb – marked with an asterisk (*) – outperform FineWeb2. For comparison, we also include EPFML High. Here, 1,000 training steps equates to approximately 2.1 billion tokens. Individual benchmark results are provided in App.  D.1.
@@
We trained and evaluated our filtered Common Crawl (Filtered CC), synthetic, and our FineWeb2 High bucket, and compared these to random sampling German FineWeb2 and the top 10% of German FineWeb2 documents classified by Messmer et al. [31] (EPFML High). The results are shown in Figure 1. We chose to exclude TruthfulQA from the 1B parameter model evaluations, as we found that it did not give a clear signal and suspect that this size of the model may be incapable of the task (cf. §3.3 for results of an 8B model).
@@
Figure 2 compares the performance of these models pre-trained for different datasets and training scenarios. Differently to the 1B Llama-style models, these 8B HAT models operate on words, not tokens, resulting in 1,000 training steps equating to approximately three billion words in this set-up. We confirmed that our synthetic data on average outperforms randomly-sampled FineWeb2 data in Figure 2 (centre). We also confirmed our previous finding of our filtered CC data outperforming FineWeb2 at the 1B scale in Figure 2 (right).
When pre-training LLMs, web data is often mixed from different languages and with data from alternative, curated sources [16], e.g., English and German data might be mixed with additional non-web data from books and encyclopedias. To test the performance of our single highest-performing subset, we compared our synthetic data to mix of 50% FineWeb2 and 50% high-quality curated datasets [^4] In particular: the German National Library, European Union Parliament transcripts, European Patent Office applications, German Text Archive, Project Gutenberg books, JRC-Acquis, NewsAPI, Open Legal Data, Open Web Math, peS2o, PhilPapers, Wikibooks, Wikinews, Wikpedia, Wikisource, and Wikivoyage.. Figure 2 (left) shows our synthetic data outperformed this mixture on our tested benchmarks.
Figure 2: Accuracies of 8B HAT models trained on different datasets. On average, and for all but one individual benchmark comparison, Aleph-Alpha-GermanWeb outperforms FineWeb2. (Left) The HAT model was trained for 25,000 training steps on an English-language web-derived dataset, equating to $\sim$ 75 billion English words. Afterwards the model was further trained for 20,000 steps ($\sim$ 60 billion words) with German language data, either random FineWeb2 data augmented with high quality datasets like Wikipedia (FineWeb2:HQ Curated) or our generated synthetic data (Synthetic). Both datasets amount to 60 billion German words each. (Middle) We trained for 21,000 steps ($\sim$ 63 billion words) on either random FineWeb2 data or on our synthetic German data. (Right) The FineWeb2 training run was continued to 50,$0$ steps and is compared to a training of 50,000 steps on data from our Filtered CC pipeline. Both runs each equate to $\sim$ 150 billion words.
@@ 4 Discussion and Conclusions
Data preparation: Michael Barlow, Bastian Harren, Volker Stampa, Stephan Wäldchen, Gregor     Ziegltrum
@@ Appendix D Extended results
Aleph-Alpha-GermanWeb outperforms FineWeb2 across all benchmarks. Interestingly, we find that different subsets have different strengths: our synthetic data performed particularly well on MMMLU, whereas our filtered CC and high-quality classified data performed best for HellaSwag and ARC.
@@
D.2 Compute Resources
The training of our 1B-model for 40k steps, which in total amounts to roughly $8.4\times 10^{10}$ tokens, was performed for 9 datasets: Our 5 FineWeb2 quality buckets, random Fineweb, epfml-FineWeb2, Filtered CC, and our synthetic data. Computation was done with current standards in bfloat16 format. The total compute budget was 12 hours on 64 H100 GPUs per ablation.
The 8B HAT-model was trained for 3 different experiments with 5 runs in total. First, for 50k steps, which in total amount to roughly $1.5\times 10^{11}$ words for Filtered CC and FineWeb2. Secondly, for 21k steps, around $6.3\times 10^{10}$ words for synthetic data. The 21k snapshot from the FineWeb2 training was reused for this comparison. Lastly for 25k on English and 20k on German data, equating to roughly $1.35\times 10^{11}$ words in total. Note that each word is one token. The total compute budget was 4 days on 128 H100 GPUs per ablation.